OpenStudy (anonymous):

.Graph the first six terms of a sequence where a1 = 4 and r = 2.

OpenStudy (anonymous):

@Haseeb96 @myko @mathstudent55

OpenStudy (anonymous):

11.jpg

OpenStudy (anonymous):

i get confused :c

OpenStudy (anonymous):

geometric series. Every term is obtained by multiplying the previous by some number r

OpenStudy (anonymous):

first term a_1=4 second a_2=4*r=4*2=8 and so on

OpenStudy (anonymous):

Oh, So can you help me with one more? ;o

OpenStudy (anonymous):

post other question